OpenStudy (anonymous):

Factor the trinomial below. 9x2 + 6x + 1

OpenStudy (anonymous):

3?

OpenStudy (kuoministers):

you need 2 numbers

OpenStudy (aravindg):

first of all \[\huge \color{green}{\text{WELCOME TO OPENSTUDY!}}\]

OpenStudy (anonymous):

Thanks Aravind!

OpenStudy (kuoministers):

\[\Huge{\color{Green}{\rightarrow \boxed{\mathbb{\text{WELCOME}{}}}{\color{green}{\color{blue}TO{\color{purple}O!}}}}} \]

OpenStudy (anonymous):

Sorry and Thanks you!

OpenStudy (mathmate):

Looks like a perfect square of the form (a+b)^2=a^2+2ab+b^2 if you look at the question as: (3x)^2+6x+1

OpenStudy (anonymous):

Solve the equation using the quadratic formula, equal the whole expression to 0 and proceed, then you'll obtain 2 values of x, for example, if you get x=1 and x=2, to put it as a product you must put it with the sign changed, in this example it would be (x-1)(x-2). Sometimes you can't use the quadratic formula, I suggest you learn to factor in this order: 1-Common Factor. 2-Notable pars. 3-Ruffini Rule (you should always use this in order to factor powers above 2) 4-Quadratic formula can only be applied when you have an equation in the form Ax^2 + By + c = 0 \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

Thanks